Liquid hexane
will react with gaseous oxygen
to produce gaseous carbon dioxide
and gaseous water
. Suppose 60.3 g of hexane is mixed with 53. g of oxygen. Calculate the minimum mass of hexane that could be left over by the chemical reaction. Round your answer to
significant digits.
1
Expert's answer
2018-02-13T06:23:13-0500
Solution
The balanced equation for above reaction is:
2 C 6 H 14 + 19 O 2 = 12 CO 2 + 14 H 2 O
The given mass of hexane would require:
60.3 g C 6 H 14 (1 mol C 6 H 14 /86.175 g C 6 H 14 )(19 mol O 2 /2 mol C 6 H 14 )(32 g O 2 /1 mol O 2 ) =
= 212.7 g O 2 .
Since this amount of oxygen is not present, we can conclude that oxygen is the limiting
reactant.
The mass of hexane that would react with 53. g of oxygen is:
53.0 g O 2 (1 mol O 2 /32 g O 2 )( 2 mol C 6 H 14 / 19 mol O 2 )( 86.175 g C 6 H 14 /1 mol C 6 H 14 ) =
= 15.0 g C 6 H 14 .
Therefore, the minimum mass of hexane that could be left over by the chemical reaction is:
60.3 g – 15.0 g = 45.3 g.
Answer: 45.3 g.